OpenStudy (anonymous):

What is the slope of the hypotenuses of the triangles in simplest form? @ganeshie8 @dan815 @waterineyes

OpenStudy (anonymous):

@waterineyes can you help me

OpenStudy (shinalcantara):

slope is defined by the formula: \[m = \frac{ y_2 - y_1 }{ x_2 - X_1 }\] where (x1,y1) , (x2,y2) are the points of the line. now try substitute the formula given the points =)

OpenStudy (shinalcantara):

note that the two triangles are just having the same slope as it lies on the same line

OpenStudy (anonymous):

i still don't get it

OpenStudy (shinalcantara):

let's take the smaller triangle's slope. isn't it that the hypotenuse's endpoints are (3,1) and (6,2) ? take it then as the points (x1,y1) and (x2,y2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so wats the anwer

OpenStudy (shinalcantara):

have you tried substituting the values?

OpenStudy (anonymous):

is the anwer 1/3

OpenStudy (shinalcantara):

\[m = \frac{ y_2-y_1 }{ x_2-x_1 }\] \[m = \frac{ 2-1 }{ 6-3 }\]

OpenStudy (shinalcantara):

yep.. that's the slope..

OpenStudy (anonymous):

so that's the anwer

OpenStudy (shinalcantara):

now try to get the slope of the bigger triangle with the same idea that you have about slope. this time you'll have the points as (6,2) and (18,6)

OpenStudy (shinalcantara):

though we know that you'll get the same answer since it lies on the same line, we just have to prove that our idea that they are the same is correct

OpenStudy (anonymous):

4/12 ???

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

reduce it to lowest term 4/12 is also 1/3 xD

OpenStudy (anonymous):

so 1/3 is the anwer?

OpenStudy (shinalcantara):

yep.

OpenStudy (shinalcantara):

just remember the formula for slope: \[m = \frac{ y_2-y_1 }{ x_2-x_1 }\] where m is the slope and (x1,y1) and (x2,y2) are the two points of the line